# AIDOVECL: AI-generated Dataset of Outpainted Vehicles for Eye-level Classification and Localization

### v1

@@ 4.1 Outpainted Vehicles
4.2 Object Classification and Localization
In an ablation study examining AIDOVECL’s capability to augment real datasets, we train YOLOv8 models on real images augmented by AIDOVECL, generated solely for the training split and doubling its size (one outpainted image per real image). The validation and test splits consist only of real images. This method assesses how augmentation with outpainted vehicles influences model performance on unseen real images, relative to using only the real dataset. In addition, we evaluate AIDOVECL affect on metrics in the presence of two other augmentation techniques, i.e. mixup and mosaic, as shown in Table 3. The confusion matrices in Figure 3 complements the results by demonstrating improvements over different classes.
As Table 3 shows, the augmented dataset outperforms the real one in almost all cases. The improvement in precision is more pronounced than in recall, likely because AIDOVECL does not alter the brand and model of the car, but changes the color, size, and location of it in the scene. Therefore, recall, which measures the diversity of correctly identified classes, may benefit less from the outpainting method compared with precision. However, the F1 score, which incorporates both precision and recall, shows significant improvement in all studied cases. The bounding box prediction accuracies improve in all cases as indicated by mAP50, mAP50-95, and fitness. This can be attributed to the random scaling and positioning of objects in AIDOVECL. Overall, while the Augmented case may not show significant standalone improvement over Real (mosaic) and Real (mixup), it excels when combined with these techniques because the introduction of outpainted vehicles adds unique contextual variations that blend effectively with the randomized environments created by mixup and mosaic. This combination enriches the dataset’s diversity far beyond that of the real dataset alone, enhancing the model’s ability to handle complex, varied inputs effectively.
Figure 3 provides detailed predictions for real and augmented datasets across different vehicle classes using confusion matrices. The numbers in parentheses indicate the count of vehicles, and the fractions in each cell represent the proportion of predictions for that class, with the sum of percentages in each column equaling one. Augmenting with AIDOVECL significantly enhances the representation of underrepresented classes (2-4, 6, and 7), demonstrating the effectiveness of outpainting as a method to address class imbalance. However, outpainting vehicles can lead to some confusion; for example, COUPES (0) and PICKUPS (7) are more likely to be confused with SEDANS (1) and SUVs (2), respectively. This occurs because outpainting uses a faded mask to maintain scene integrity, which may obscure features like the two doors on a COUPE or artificially extend the trunk of a PICKUP to resemble an SUV roof. Despite these few instances of increased confusion, the general trend shows improved predictions when using AIDOVECL, which also significantly reduces the occurrence of false positives.
Table 3: Classification and localization performance metrics for real datasets and those augmented by AIDOVECL, with additional evaluations using mixup and mosaic techniques for both cases.
Case Precision Recall F1 Score mAP50 mAP50-95 Fitness
Real 0.748 0.817 0.781 0.842 0.786 0.791
Augmented 0.789 0.815 0.802 0.873 0.817 0.823
Real (mosaic) 0.813 0.791 0.802 0.874 0.813 0.819
Augmented (mosaic) 0.859 0.850 0.854 0.902 0.857 0.862
Real (mixup) 0.793 0.803 0.798 0.853 0.783 0.790
Augmented (mixup) 0.874 0.827 0.850 0.910 0.851 0.857
Figure 3: Confusion matrices for the real and augmented (with AIDOVECL) datasets using mixup and mosaic augmentations.
@@ 5 Conclusion
The AIDOVECL dataset represents a significant advancement in the field of machine learning for vehicle classification and localization. It addresses the critical issue of annotated data scarcity in desired classes and perspectives, and mitigates the problem of unlabeled background objects in public datasets. By incorporating generative AI techniques—specifically prompt-guided outpainting—AIDOVECL not only enriches the diversity of eye-level vehicle images but also effectively simulates realistic urban traffic scenarios. This approach is crucial for training algorithms capable of performing accurately in dynamic real-world environments. AIDOVECL’s innovative self-annotating method, which potentially forgoes the need for manual annotation, offers a substantial improvement over traditional dataset curation methods. It can augment real datasets with class imbalances, thus enhancing the adaptability and performance of machine learning models. This dataset meets immediate training needs and fosters future advancements in computer vision technologies for applications ranging from autonomous driving to urban planning.
Limitations.
A primary limitation is that we currently cannot outpaint multiple vehicles within a single image, as our employed inpainting model fails to generate reasonable and coherent scenes with two or more desired vehicles. This challenge stems not only from limitations of the inpainting model but also from the difficulty in selecting correlated seed images that realistically might be observed together. This limitation restricts our dataset’s ability to provide rich occlusion cases for objects, which are crucial for training robust detection systems. Additionally, our reliance on pretrained detection and inpainting models for generating seed images and outpainting them presents another limitation. If the objects deviate significantly from those used during model training, they may not be detected, leading to their omission from the pool of seed images, or the outpainting may appear unrealistic. Moreover, there is a need for more images, as seed images must be selected from clean images that meet specific criteria. Therefore, we still need more real data to generate more diverse outpainted images. Addressing these limitations will be essential for enhancing the functionality and applicability of our dataset, ultimately leading to more accurate and versatile object detection models.
Future Works.
To address the existing limitations, enhancing inpainting models with spatial transformer networks or advanced attention mechanisms can significantly improve scene coherence, particularly in handling complex scenarios involving multiple objects. These technologies enable the model to focus selectively on different parts of an image and adjust the spatial relationships between objects, leading to more realistic and dynamically composed scenes. Additionally, advancing data curation methods is crucial. Employing machine learning algorithms that can analyze and predict plausible object-background pairings based on contextual and spatial data could revolutionize how seed images are selected and combined. This approach not only enhances the realism of the generated images but also ensures that the synthetic data better mimics real-world variability, thereby improving the training and performance of object detection models.
@@
In this research, we introduce a method for generating training datasets for vehicle detection and classification using outpainting techniques. The potential benefits of our work include enhanced accuracy in automated vehicle detection, which could improve urban planning, traffic management, and autonomous driving, contributing to safer and more efficient urban environments. While our dataset is carefully curated to focus on vehicle features that support these applications, we acknowledge that any form of data capture involving vehicles may contain elements that could raise privacy concerns. We emphasize the importance of ethical guidelines and responsible use of technologies to mitigate potential risks associated with data misuse.
@@
The authors acknowledge support from the University of Illinois NCSA Faculty Fellowship Program, the Zhejiang University and University of Illinois Dynamic Research Enterprise for Multidisciplinary Engineering Sciences (DREMES), and the University of Illinois Discovery Partners Institute (DPI). This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its NCSA  (Gropp et al., 2023).

### v2

@@ 4.2 Outpainted Vehicles
Figure 4: KDE plots for visual quality scores (a–c) and semantic similarities of prompt and captions (d–f). Larger values generally indicate better quality and semantics, except for BRISQUE where smaller values are better.
4.3 Overall Detection Performance Across Models
We conducted an ablation study that spans five modern object‑detection architectures (FCOS, RetinaNet, SSD‑Lite, FasterRCNN, and YOLOv8). For every model we evaluate four train/test configurations—Real/Real, Aug./Real, Real/Aug., and Aug./Aug.—to isolate the benefit of introducing synthetic data during training and the robustness to distribution shifts at inference. Here, augmentation signifies that each real image is accompanied with a synthetic counterpart from AIDOVECL; the generation pipeline surpasses a 91% success rate against our quality criteria after up to 100 iterations per image. Within every configuration we further vary on‑the‑fly augmentations—namely the mixup ratio and mosaic probability—to probe how aggressive blend‑and‑compose policies interact with the underlying data composition. Six standard metrics (Precision, Recall, F1, mAP50, mAP50‑95, and Fitness) are reported for all combinations.
Part of the results in Tables 3, 4, and 5 compare a baseline scenario—where each detector is both trained and evaluated on real images—to a scenario where the training dataset is augmented with AIDOVECL, while evaluation remains solely on real images. That is, we first focus on "Real/Real" and "Aug./Real" results. Table 3 demonstrates that adding synthetic AIDOVECL images improves almost all metrics when no additional on-the-fly augmentations are applied. Improvements of up to 10% are noted across each model. Increasing the blending factor using Mixup (mixup=0.5) as shown in Table 4, however, does not lead to consistent or significant improvements when training data is augmented. This suggests that pixel-level blending from Mixup effectively reduces overfitting. Consequently, if inference data closely matches training distribution, substantial benefits from AIDOVECL may be limited when using Mixup simultaneously. On the other hand, Table 5 shows that when only Mosaic augmentation is applied with a probability of 0.5, there is an improvement of up to 10%. This indicates that Mosaic augmentation alone does not diminish the advantages offered by our augmentation. Therefore, when inference remains within the distribution, AIDOVECL augmentation provides clear performance benefits for most cases.
Comparing the results for "Aug./Aug." and "Real/Aug." highlights the importance of augmenting the training data with AIDOVECL. Across all models, configurations, and metrics presented in Tables 3, 4, and 5, "Real/Aug." consistently underperforms compared to "Aug./Aug.". Although augmenting with AIDOVECL does not introduce entirely new vehicle objects, but instead rescales and relocates existing vehicles into new backgrounds, it remains crucial for the model to encounter these backgrounds during training to mitigate out-of-distribution inference issues. As demonstrated in all cases across the tables, "Aug./Aug." significantly outperforms "Real/Aug." by up to approximately 40%, indicating that augmenting training data with AIDOVECL effectively prepares the models for encountering vehicles in contexts different from those present in the real training dataset.
Our comprehensive ablation study across multiple detection architectures clearly demonstrates the effectiveness and necessity of incorporating AIDOVECL-generated synthetic data during training. Synthetic augmentation notably enhances performance, with improvements of up to 10%. Importantly, synthetic augmentation substantially boosts robustness against distribution shifts at inference, achieving performance gains of up to approximately 40% when evaluation contexts differ from training. Thus, augmenting real datasets with AIDOVECL significantly strengthens model reliability and generalization in varied deployment scenarios.
Model Train/Test Metrics
Precision Recall F1 Score mAP50 mAP50-95 Fitness
FCOS Real / Real 0.63 0.94 0.75 0.63 0.42 0.44
Aug. / Real 0.65 0.97 0.78 0.65 0.44 0.46
Real / Aug. 0.56 0.90 0.69 0.56 0.41 0.42
Aug. / Aug. 0.69 0.98 0.81 0.69 0.53 0.55
RetinaNet Real / Real 0.60 0.79 0.68 0.60 0.45 0.46
Aug. / Real 0.61 0.81 0.70 0.61 0.45 0.47
Real / Aug. 0.53 0.74 0.62 0.53 0.40 0.42
Aug. / Aug. 0.67 0.84 0.74 0.67 0.54 0.55
SSD Real / Real 0.74 0.83 0.78 0.74 0.64 0.65
Aug. / Real 0.77 0.85 0.81 0.77 0.65 0.67
Real / Aug. 0.55 0.70 0.62 0.55 0.45 0.46
Aug. / Aug. 0.74 0.84 0.79 0.74 0.62 0.64
FasterRCNN Real / Real 0.76 0.92 0.83 0.76 0.57 0.59
Aug. / Real 0.77 0.94 0.84 0.77 0.58 0.60
Real / Aug. 0.66 0.86 0.75 0.66 0.50 0.52
Aug. / Aug. 0.78 0.93 0.85 0.78 0.63 0.65
YOLOv8 Real / Real 0.81 0.83 0.82 0.86 0.81 0.81
Aug. / Real 0.88 0.86 0.87 0.91 0.87 0.87
Real / Aug. 0.79 0.79 0.79 0.82 0.76 0.77
Aug. / Aug. 0.87 0.85 0.86 0.90 0.86 0.87
Table 3: Evaluation metrics with on-the-fly data augmentation during training configured as mixup=0, mosaic=0.
Model Train/Test Metrics
Precision Recall F1 Score mAP50 mAP50-95 Fitness
FCOS Real / Real 0.71 0.98 0.83 0.71 0.47 0.50
Aug. / Real 0.70 0.97 0.82 0.70 0.49 0.51
Real / Aug. 0.65 0.98 0.78 0.65 0.46 0.48
Aug. / Aug. 0.76 0.98 0.86 0.76 0.61 0.63
RetinaNet Real / Real 0.73 0.91 0.81 0.73 0.56 0.58
Aug. / Real 0.71 0.88 0.79 0.71 0.52 0.54
Real / Aug. 0.62 0.83 0.71 0.62 0.48 0.49
Aug. / Aug. 0.76 0.90 0.82 0.76 0.61 0.63
SSD Real / Real 0.74 0.88 0.80 0.74 0.61 0.62
Aug. / Real 0.71 0.83 0.77 0.71 0.59 0.61
Real / Aug. 0.54 0.73 0.62 0.54 0.43 0.44
Aug. / Aug. 0.70 0.83 0.76 0.70 0.58 0.59
FasterRCNN Real / Real 0.79 0.98 0.88 0.79 0.54 0.57
Aug. / Real 0.80 0.98 0.88 0.80 0.57 0.60
Real / Aug. 0.69 0.93 0.79 0.69 0.50 0.52
Aug. / Aug. 0.81 0.98 0.88 0.81 0.62 0.64
YOLOv8 Real / Real 0.82 0.83 0.83 0.87 0.82 0.83
Aug. / Real 0.86 0.86 0.86 0.90 0.86 0.87
Real / Aug. 0.80 0.81 0.80 0.84 0.79 0.79
Aug. / Aug. 0.86 0.85 0.86 0.90 0.86 0.86
Table 4: Evaluation metrics with on-the-fly data augmentation during training configured as mixup=0.5, mosaic=0.
Model Train/Test Metrics
Precision Recall F1 Score mAP50 mAP50-95 Fitness
FCOS Real / Real 0.61 0.94 0.74 0.61 0.41 0.43
Aug. / Real 0.67 0.95 0.78 0.67 0.47 0.49
Real / Aug. 0.66 0.95 0.78 0.66 0.50 0.52
Aug. / Aug. 0.74 0.95 0.83 0.74 0.60 0.61
RetinaNet Real / Real 0.70 0.84 0.76 0.70 0.54 0.56
Aug. / Real 0.74 0.90 0.81 0.74 0.56 0.58
Real / Aug. 0.72 0.86 0.78 0.72 0.59 0.60
Aug. / Aug. 0.78 0.92 0.84 0.78 0.65 0.66
SSD Real / Real 0.71 0.81 0.75 0.71 0.61 0.62
Aug. / Real 0.72 0.83 0.77 0.72 0.60 0.62
Real / Aug. 0.68 0.80 0.74 0.68 0.57 0.58
Aug. / Aug. 0.72 0.83 0.77 0.72 0.62 0.63
FasterRCNN Real / Real 0.68 0.92 0.78 0.68 0.47 0.49
Aug. / Real 0.69 0.96 0.80 0.69 0.47 0.49
Real / Aug. 0.68 0.90 0.77 0.68 0.51 0.53
Aug. / Aug. 0.75 0.95 0.84 0.75 0.58 0.59
YOLOv8 Real / Real 0.83 0.81 0.82 0.86 0.81 0.82
Aug. / Real 0.84 0.87 0.85 0.90 0.86 0.86
Real / Aug. 0.81 0.81 0.81 0.85 0.79 0.80
Aug. / Aug. 0.85 0.85 0.85 0.89 0.85 0.86
Table 5: Evaluation metrics with on-the-fly data augmentation during training configured as mixup=0, mosaic=0.5.
4.4 Class-wise Performance of YOLOv8 Detection
In this section, we examine detection performance at the class level, unlike the previous section which focused on overall dataset metrics. We use the YOLOv8 model, known for its strong balance between accuracy and computational efficiency, to analyze confusion matrices and visual examples of false positives, false negatives, and class confusions. This class-wise view helps us better understand where the model struggles and how augmentation with AIDOVECL leads to improvements, particularly for rare, visually similar, or previously hard-to-detect classes. By illustrating representative samples, we analyze which mismatches disappear, arise, or endure when the training data is augmented with AIDOVECL—hereafter referred to as resolved, emergent, and persistent mismatches.
Confusion Matrices.
Figures 5 and 6 show confusion matrices for the YOLOv8 model tested on real and augmented data, respectively. In each case, the model is trained on both real and augmented datasets under various on-the-fly augmentation configurations. Overall, confusion, false negatives, and false positives generally decrease across classes when the model is trained on augmented data compared to real data, and this improvement becomes even more pronounced when testing is also conducted on augmented data. Exposure to vehicles in varying sizes and contextual backgrounds during training enhances the model’s ability to detect them during testing, as it learns to generalize across spatial scales and diverse scenes. In over 80% of class-configuration combinations, the use of AIDOVECL leads to an increase in true positive counts. The most substantial gain—approximately a 50% increase in true positives—is observed for the underrepresented van class, likely because such classes have more room for performance improvement compared to already well-represented ones. Notable reductions in confusion include fewer misclassifications of vans as minivans or minibuses, coupes and SUVs as sedans, and a decrease in mutual confusion between minibuses and buses, with a general decline in mutual confusion among similar vehicle types. For other classes, improvement or deterioration in confusion is relatively minor. AIDOVECL helps to reduce overall false negatives and false positives in almost all ablation cases, because it increases the frequency of object appearances in diverse contexts, making the model less likely to miss true objects and less prone to trigger detections on background patterns.
Figure 5: Confusion matrices of the YOLOv8 model tested on the real dataset under different augmentation settings, including dataset augmentation with AIDOVECL and on-the-fly MixUp/Mosaic.
Figure 6: Confusion matrices of the YOLOv8 model tested on the augmented dataset under different augmentation settings, including dataset augmentation with AIDOVECL and on-the-fly MixUp/Mosaic.
Confused Instances.
Confused instances are those in which the predicted bounding box closely matches the ground truth location, but the predicted class label does not. Figure 7 presents examples of such instances across the previously defined categories—resolved, emergent, and persistent—after training with the AIDOVECL-augmented dataset and testing on real data without MixUp or Mosaic augmentations. Notably, the number of emergent confused instances is roughly half that of the resolved ones, which aligns with the trends observed in the confusion matrices. Interestingly, at least three emergent examples—specifically, f1, e2 and e3—demonstrate more accurate bounding box predictions than the corresponding ground truth annotations, despite the class confusion. This highlights the model’s ability—when trained on augmented data—to localize objects precisely even when class prediction becomes challenging.
False Negative Instances.
Occasionally, no prediction is made for a ground truth object, or the predicted bounding box—though class-correct—falls below the IoU threshold; such cases are considered false negatives. As shown in Figure 8, these instances are more often resolved than newly introduced when training with augmented data. Most false negatives involve vehicles that are partially occluded, visually faded due to distance or weather, or have unusual body shapes not seen during training. In some persistent cases, the YOLOv8 model correctly detects vehicles that are missing from the ground truth due to annotation constraints—each image is labeled with only the largest bounding box, omitting smaller vehicles as in a5, d5, b8, and d8. Though rare, such examples demonstrate the model’s ability to detect background vehicles missed by PyTorch detectors, particularly when trained on augmented data (e.g., compare a5 vs. d5, where Torch detected a person as a vehicle by mistake). In other cases—often involving articulated trucks—YOLOv8 even outperforms the ground truth annotations by Torch, so the resulting false negative count underestimates actual performance.
False Positive Instances.
The YOLOv8 model may also produce false positives by detecting vehicles for which no corresponding ground truth exists. Such cases often stem from inaccurate, partial, or missing annotations, and more rarely from background elements—such as buildings—being mistakenly detected as vehicles. As shown in the persistent samples of Figure 9, smaller vehicles (typically buses) omitted from annotations, or partially labeled trucks, are successfully detected by the YOLOv8 model. Notably, training with augmented data is more likely to yield accurate bounding boxes in such cases (e.g., compare a7 vs. d7 and b7 vs. e7). Instances of accurate detection of partially annotated vehicles may appear in both false negative and false positive categories because the predicted box, though correct, may not sufficiently overlap with the ground truth due to the annotation’s smaller size or limited extent. When the IoU falls below the matching threshold, this results in a false negative from the ground truth’s perspective and a false positive from the model’s, as neither considers the other an acceptable match.
Figure 7: Confused instances, where the ground truth is correctly localized but misclassified by the prediction, shown as GT (solid): [class] (confused) and Pred#[n]: [class] (confidence, IoU).
Figure 8: False negative instances, where the ground truth is either not detected or only partially matched with an IoU below the threshold, shown as GT (solid): [class] and Pred#[n] FP: [class] (confidence).
Figure 9: False positive instances, where a prediction corresponds to no or only partially matched ground truth, shown as GT (solid): (matched/unmatched/confused) [class] and Pred#[n] (FP/TP/Conf): [class] (confidence, IoU).
5 Limitations
Our reliance on pretrained detection and inpainting models introduces certain limitations, particularly when dealing with non-standard or atypical vehicles that may not be recognized during seed extraction or may result in unrealistic outpainting. These limitations are most relevant when the goal is to include classes not well represented in the pretrained models’ training data. However, this does not undermine our claims, as our focus—reflected throughout the paper and in the use of the term vehicle—has been on standard, commonly encountered types. Importantly, the proposed pipeline remains flexible: additional seed images or new object classes can be easily incorporated to enhance robustness within existing categories or to support training models that target broader or more specialized vehicle types.
When trying to outpaint multiple objects on the same canvas, the model faces several challenges. With only general scene cues (such as location or timing), it must guess each object’s placement and relation to the scene from mask geometry and surrounding pixels. In the LDM architecture, the entire masked area is encoded and denoised as a single spatially coherent latent feature map, with no mechanism to treat disconnected regions independently. If these masked areas are handled in a single pass, the model processes them as one editing region, leading to cross-attention competition—where one object dominates and others may suffer from omissions or partial presence (i.e., objects that are faint or blended into the background, making them no longer clearly distinct). This one-pass complexity forces the model to resolve all spatial, style, and detail constraints simultaneously. At the same time, it must maintain global coherence—such as consistent lighting, shadows, and perspective—while ensuring smooth background blending at multiple transition boundaries, making errors more likely than when outpainting a single object. The problem is further amplified if the pretrained model has rarely or never seen such objects co-occurring in its training data. Although not a replacement for real multi-object training data, techniques like MixUp and Mosaic can help fill the gap by encouraging generalization to more complex, composite, or partially occluded scenes. Moreover, this limitation becomes less critical when AIDOVECL is seen as part of a broader augmentation pipeline rather than a standalone dataset.
Despite strict visual and perceptual vetting using quality assessment tools, manual inspection reveals that approximately 5.5% of the outpainted images still exhibit various artifacts. See the affected outpainted images in the validation split of the data in Figure 10. Heavy vehicles such as trucks and buses may suffer from non-smooth transition boundaries, patchwork effects, or irrelevant backgrounds, although these artifacts are rare. The main culprit is that the pretrained LDM was not trained specifically on vehicle-centric scenes, especially not on less frequent categories like trucks and buses. Unwanted vehicles, which are not automatically annotated, may still appear despite negative prompts including different vehicle types, likely due to their co-occurrence with other elements in the training data of the LDM. Also, we do not have any mechanism to rule out distorted roads or seasonal mismatches between regions of the image. Defective vehicles may also appear, as detection models for seed vehicle extraction may not fully capture vehicles—typically trailers. These rare cases remain usable for the downstream task of detection and are retained to benchmark our augmentation pipeline, as they err on the side of caution by potentially underestimating our performance. However, they are excluded from the curated dataset to ensure full realism and aesthetic quality, as noted in Data Availability and User Framework.
Figure 10: Artifacts are found in 5.5% of outpainted vehicles, despite strict quality checks. These cases are manually excluded from the curated dataset for aesthetic and realism.
6 Future Works
Active learning offers a promising direction for facilitating efficient seed pool construction in AIDOVECL by prioritizing which unlabeled vehicle crops—initially unclassified and manually assigned to custom categories—should be classified first. In a cold-start setting, diversity-first selection through core-set methods can ensure that the earliest labeled vehicle crops are representative of the broader data distribution (sener2017active). Task-agnostic latent-space selection, as in variational adversarial active learning, can further identify underrepresented or heterogeneous vehicle types before a classifier is trained (sinha2019variational). Once a small classifier exists, uncertainty-based querying using Bayesian methods can target ambiguous or confusing vehicle crops (gal2017deep), while batch acquisition strategies such as BADGE can balance informativeness and diversity when expanding seeds in groups (ash2019deep). In addition, loss prediction modules can highlight samples expected to be difficult for the classifier (yoo2019learning), ensuring annotation effort is concentrated where it yields the most improvement. Finally, self-training approaches can automatically assign labels to high-confidence samples while leaving manual effort for low-confidence or rare vehicle types (xie2020self). Together, these approaches could reduce manual labeling costs and accelerate seed pool growth, thereby strengthening the foundations of AIDOVECL.
To overcome the limitations of single-pass multi-object outpainting—where disconnected regions compete for attention and global coherence is difficult to maintain—several promising strategies can be drawn from recent advances in conditional and controllable diffusion models: One approach is to guide generation with the original image (anchor view) and location-aware features (relative positional embeddings) (zhang2024continuous). Alternatively, a grounded outpainting model inspired by GLIGEN (li2023gligen) could condition on inputs defining the canvas layout to preserve object positions while generating coherent context. Structural inputs such as depth or edge maps, used in ControlNet (zhang2023adding), can guide outpainted regions around multiple objects to follow scene geometry and blend naturally into the environment. MultiDiffusion (bar2023multidiffusion) can be adapted for multi-object outpainting by coordinating generation across background and object-adjacent regions so that lighting, perspective, and style remain consistent. Composable diffusion methods (liu2022compositional) can improve multi-object handling by combining independent condition scores to reduce concept interference. Attend-and-Excite (xu2018attngan), while originally developed for text-to-image generation, can inspire methods that keep attention balanced across multiple objects, helping preserve coherence in their expanded surroundings. All these methods require fine-tuning of existing diffusion models to incorporate the additional conditioning and control mechanisms they rely on. Moreover, advancing data curation with ML to predict plausible object–background pairings can further enhance realism and ensure synthetic data better reflects real-world variability.
Future work may explore strategies to further reduce the occurrence of outpainting artifacts. Fine-tuning the pretrained LDM on vehicle-centric datasets—including underrepresented classes such as trucks and buses—could improve generation fidelity for those categories. Exploring state-of-the-art scoring mechanisms, such as artifact detectors, might help identify and discard images with non-smooth transitions, patchwork effects, or irrelevant backgrounds more effectively. Additionally, refining prompt engineering techniques or applying context-aware negative prompting may help suppress unwanted vehicles, distorted roads, seasonal mismatches, and ambiguous objects more reliably. Overall, while most generated images are usable for detection tasks, improving visual realism and scene consistency remains an important direction.
@@ 7 Conclusion
The AIDOVECL dataset represents a significant advancement in the field of machine learning for vehicle classification and localization. It addresses the critical issue of annotated data scarcity in desired classes and perspectives, and mitigates the problem of unlabeled background objects in public datasets. By incorporating generative AI techniques—specifically prompt-guided outpainting—AIDOVECL not only enriches the diversity of eye-level vehicle images but also effectively simulates realistic urban traffic scenarios. This approach is crucial for training algorithms capable of performing accurately in dynamic real-world environments. AIDOVECL’s automatic annotation, which potentially forgoes the need for manual annotation, offers a substantial improvement as demonstrated across various detection models, with or without on-the-fly augmentation methods. It can augment real datasets with class imbalances, thus enhancing the adaptability and performance of machine learning models. The dataset and its associated pipeline meet immediate training needs and fosters future advancements in computer vision technologies for applications ranging from autonomous driving to urban planning.
Dataset Availability and User Framework
The AIDOVECL resource is released both as (i) a dataset comprising curated and benchmark versions and (ii) a modular framework enabling end-users to generate their own augmented datasets. The dataset is openly hosted on Hugging Face at https://huggingface.co/datasets/amir-kazemi/aidovecl, accompanied by a dataset card documenting the structure, splits, and annotation format. The dataset is distributed under a CC-BY-4.0 license. The generation pipeline, with full source code and usage instructions, is publicly available on GitHub at https://github.com/amir-kazemi/aidovecl under the MIT license.
We provide both a curated dataset—combining real and AI-generated samples in a conventional split ratio for computer vision tasks, with rare cases containing artifacts excluded through manual inspection—and a research archive of benchmark datasets for reproducing the paper’s results, which adopt a higher test split ratio to enable more rigorous evaluation. The latter is organized as follows: Raw unvetted images are provided under real_raw (with class-wise subfolders for buses, coupes, minibuses, minivans, pickups, sedans, SUVs, trucks, and vans). From these, vetted seed images are extracted and stored under real, which contains images/, labels/, and a real.yaml configuration file for YOLO-style training. Augmented data produced by AIDOVECL is placed under augmented along with real ones in a corresponding directory, with a parallel organization of images/, labels/, and augmented.yaml.
The pipeline is intended not only as a static dataset release but also as a tool for researchers to generate additional augmented data tailored to their own target domains. Users are encouraged to apply the framework to broaden vehicle classes, adapt to novel environments, or scale dataset size as needed. In such cases, they are required to cite the present AIDOVECL work as the originating framework. While we do not plan to expand the official dataset with additional vehicle categories at this stage, we will strive to address reported bugs and ensure that the repositories remain usable and well-documented. In this way, the dataset and framework are both stable and extensible, supporting community-driven evolution.
@@
In this research, we introduce a method for generating training datasets for vehicle detection and classification using outpainting techniques. The potential benefits of our work include enhanced accuracy in automated vehicle detection, which could improve urban planning, traffic management, and autonomous driving, contributing to safer and more efficient urban environments. While our dataset is carefully curated to focus on vehicle features that support these applications, we acknowledge that any form of data capture involving vehicles may contain elements that could raise privacy concerns. We emphasize the importance of ethical guidelines and responsible use of technologies to mitigate potential risks associated with data misuse. In particular, AIDOVECL is intended primarily for research and educational purposes, not for surveillance or applications that could infringe on individual rights (although because AIDOVECL only includes labels for vehicle class, we believe it to be of minimal utility for such applications). Users of the dataset and framework are expected to comply with relevant legal regulations and institutional review requirements, and to exercise caution when extending the pipeline to new domains. To support transparency, the dataset (CC-BY-4.0) and code (MIT) licenses are explicitly provided in their respective repositories.
@@
The authors acknowledge support from the University of Illinois NCSA Faculty Fellowship Program, the Zhejiang University and University of Illinois Dynamic Research Enterprise for Multidisciplinary Engineering Sciences (DREMES), and the University of Illinois Discovery Partners Institute (DPI). This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its NCSA (Gropp2023Delta).